OpenStudy (anonymous):

A rectangle has an area of 16 square feet. Its length and width are whole numbers. What is its maximum perimeter? 32 feet 64 feet 34 feet 16 feet

OpenStudy (anonymous):

16 = xy x = 4, y = 4 P = 2x + 2y P = 8+8 P = 16

OpenStudy (anonymous):

64

OpenStudy (anonymous):

I think the function is maximized at (4, 4)